# Exponentiations of ultrafilters

### v4

@@ 2   Exponentiations of ultrafilters
First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual method of applying Ellis’ Theorem (see [5, Theorem 2.5] for details). Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”. Our starting observation is the following:
@@
Assume that there exist $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ such that $p=p\odot E_{1}(q,p)$. Then $E_{1}(q,p)$ is $E_{1}$-idempotent.
Analogously, if there exist $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ such that $p=E_{2}(p,q)\odot p$ then $E_{2}(p,q)$ is $E_{2}$-idempotent.
@@
By [7, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\ldots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
@@
Unfortunately, we have not been able to prove the above conjecture; we can, however, show a partial result towards its solution. The basic observation is the following: by Proposition 2.10, if $p$ is $E_{2}$-idempotent, we have that
@@
By contrast, let $p$ be such an ultrafilter. Let $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n$ to the sum of the exponents in its prime factorization. In particular, $\hat{\Omega}(p)=\hat{\Omega}(E_{2}(p,p))=p\odot\hat{\Omega}(p)$ and $\hat{\Omega}(p)=\hat{\Omega}(p\odot p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, therefore $\hat{\Omega}(p)$ is additively idempotent and $p\odot\hat{\Omega}(p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$. If $\hat{\Omega}(p)\in\mathbb{N}$, then by additive idempotency we would get that $\hat{\Omega}(p)=0$, from which $p=1$, which is a contradiction as $p\in\beta\mathbb{N}\setminus\mathbb{N}$. So $\hat{\Omega}(p)$ is additively idempotent and nonprincipal, hence $n\mathbb{N}\in\hat{\Omega}(p)$ for all $n\in\mathbb{N}$. Thus we can apply [5, Theorem 13.14] to deduce that $p\odot\hat{\Omega}(p)\neq\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, reaching the desired contradiction. ∎
@@ 3   Open questions
For one ultrafilter, the following question, for which we conjecture a negative answer, seems relevant:
@@
Does it exist $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $E_{1}(p,p)=E_{2}(p,p)$?

### v5

@@ 2   Exponentiations of ultrafilters
First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual method of applying Ellis’ Theorem (see [5, Theorem 2.5] for details). Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only principal $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”. Our starting observation is the following:
@@
Assume that there exist $p,q\in\beta\mathbb{N}\setminus\{1\}$ such that $p=p\odot E_{1}(q,p)$. Then $E_{1}(q,p)$ is $E_{1}$-idempotent.
Analogously, if there exist $p,q\in\beta\mathbb{N}\setminus\{1\}$ such that $p=E_{2}(p,q)\odot p$ then $E_{2}(p,q)$ is $E_{2}$-idempotent.
@@
By [7, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\ldots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b>1$.
@@
As expected, a positive answer to the above conjecture would have several interesting combinatorial consequences: for example, any set belonging to such an ultrafilter would be exponentially-IP, in the following sense.
Theorem 2.31.
Let $p\in\beta\mathbb{N}$ such that $p=E_{2}(p,p)$. For each $A\in p$, there is a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ in $\mathbb{N}$ such that $\{x_{n}\mid n\in\mathbb{N}\}\cup\{x_{n+1}^{y}\mid n\in\mathbb{N}\ \text{and}\ y\in FP(\langle x_{t}\rangle_{t=1}^{n})\}\subseteq A$.
Proof.
For $A\subseteq\mathbb{N}$ and $x\in\mathbb{N}$, let $D(A,x)=\{y\in\mathbb{N}\mid y^{x}\in A\}$; for each $A\in p$, $\{x\in\mathbb{N}\mid D(A,x)\in p\}\in p$ by $E_{2}$-idempotency of $p$.
Let $A\in p$, let $A_{1}=A$, and let $B_{1}=\{x\in\mathbb{N}\mid D(A_{1},x)\in p\}$. Pick $x_{1}\in A_{1}\cap B_{1}$ and let $A_{2}=A_{1}\cap B_{1}\cap D(A_{1},x_{1})$.
Now let $n\in\mathbb{N}\setminus\{1\}$ and assume that $A_{n}\in p$. Let $B_{n}=\{x\in\mathbb{N}\mid D(A_{n},x)\in p\}$, pick $x_{n}\in A_{n}\cap B_{n}$, and let $A_{n+1}=A_{n}\cap B_{n}\cap D(A_{n},x_{n})$.
Having chosen $\langle x_{n}\rangle_{n=1}^{\infty}$, it is immediate that $\{x_{n}\mid n\in\mathbb{N}\}\subseteq A$. Now let $n\in\mathbb{N}$. We show by induction on $|F|$ that for $\emptyset\neq F\subseteq\{1,2,\dots,n\}$, if $m=\min F$ and $y=\prod_{t\in F}x_{t}$, then $x_{n+1}^{y}\in A_{m}$. If $F=\{m\}$, then $x_{n+1}\in A_{m+1}\subseteq D(A_{m},x_{m})$ so $x_{n+1}^{x_{m}}\in A_{m}$. Now assume that $|F|>1$, let $G=F\setminus\{m\}$, let $r=\min G$, and let $z=\prod_{t\in G}x_{t}$. Then $x_{n+1}^{z}\in A_{r}\subseteq A_{m+1}\subseteq D(A_{m},x_{m})$ so $x_{n+1}^{zx_{m}}=\left(x_{n+1}^{z}\right)^{x_{m}}\in A_{m}$. ∎
Unfortunately, we have not been able to prove Conjecture 2.30; we can, however, show a partial result towards its solution. The basic observation is the following: by Proposition 2.13, if $p$ is $E_{2}$-idempotent, we have that
@@
By contrast, let $p$ be such an ultrafilter. Let $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n$ to the sum of the exponents in its prime factorization. In particular, $\widehat{\Omega}(p)=\widehat{\Omega}(E_{2}(p,p))=p\odot\widehat{\Omega}(p)$ and $\widehat{\Omega}(p)=\widehat{\Omega}(p\odot p)=\widehat{\Omega}(p)\oplus\widehat{\Omega}(p)$, therefore $\widehat{\Omega}(p)$ is additively idempotent and $p\odot\widehat{\Omega}(p)=\widehat{\Omega}(p)\oplus\widehat{\Omega}(p)$. If $\widehat{\Omega}(p)\in\mathbb{N}$, then by additive idempotency we would get that $\widehat{\Omega}(p)=0$, from which $p=1$, which is a contradiction as $p\in\beta\mathbb{N}\setminus\mathbb{N}$. So $\widehat{\Omega}(p)$ is additively idempotent and nonprincipal, hence $n\mathbb{N}\in\widehat{\Omega}(p)$ for all $n\in\mathbb{N}$. Thus we can apply Theorem 2.8 to deduce that $p\odot\widehat{\Omega}(p)\neq\widehat{\Omega}(p)\oplus\widehat{\Omega}(p)$, reaching the desired contradiction. ∎
@@ 3   Open questions
One particular case of equality was discussed in Example 2.15, from which the following question arises:
@@
For which $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$ does the equality $E_{1}(p,q)=E_{2}(p,q)$ hold? Does it exist $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $E_{1}(p,p)=E_{2}(p,p)$?
@@
Acknowledgements: We thank the referee for his/her careful reading of the paper, which led to several improvements; in particular, we are thankful for the suggestion of isolating Lemma 2.12 to simplify the proofs of Proposition 2.13.(2)-(3), for proposing an explicit proof of the properties of $\widehat{H}$ used in Case 2 of Proposition 2.16, and for indicating Theorem 2.31 as a striking example of the eventual consequences of the existence of $E_{2}$-idempotents.